# Plan Then Retrieve: Reinforcement Learning-Guided Complex Reasoning over Knowledge Graphs

### v2

@@ Acknowledgements.
- Shao et al. (2024) Zhihong Shao, Peiyi Wang, Qihao Zhu, Runxin Xu, Junxiao Song, Xiao Bi, Haowei Zhang, Mingchuan Zhang, YK Li, Yang Wu, et al. 2024. Deepseekmath: Pushing the limits of mathematical reasoning in open language models. arXiv preprint arXiv:2402.03300 (2024).
@@
- Shi et al. (2025) Yaorui Shi, Sihang Li, Chang Wu, Zhiyuan Liu, Junfeng Fang, Hengxing Cai, An Zhang, and Xiang Wang. 2025. Search and Refine During Think: Autonomous Retrieval-Augmented Reasoning of LLMs. arXiv preprint arXiv:2505.11277 (2025).
@@
- Sun et al. (2023) Jiashuo Sun, Chengjin Xu, Lumingyuan Tang, Saizhuo Wang, Chen Lin, Yeyun Gong, Lionel M Ni, Heung-Yeung Shum, and Jian Guo. 2023. Think-on-graph: Deep and responsible reasoning of large language model on knowledge graph. arXiv preprint arXiv:2307.07697 (2023).
@@
- Talmor and Berant (2018) Alon Talmor and Jonathan Berant. 2018. The web as a knowledge-base for answering complex questions. arXiv preprint arXiv:1803.06643 (2018).
@@
- Tan et al. (2025) Xingyu Tan, Xiaoyang Wang, Qing Liu, Xiwei Xu, Xin Yuan, and Wenjie Zhang. 2025. Paths-over-graph: Knowledge graph empowered large language model reasoning. In Proceedings of the ACM on Web Conference 2025. 3505–3522.
- Team et al. (2024) Qwen Team et al. 2024. Qwen2 technical report. arXiv preprint arXiv:2407.10671 2 (2024), 3.
@@
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems 35 (2022), 24824–24837.
@@
- Xiong et al. (2017) Chenyan Xiong, Russell Power, and Jamie Callan. 2017. Explicit semantic ranking for academic search via knowledge graph embedding. In Proceedings of the 26th international conference on world wide web. 1271–1279.
@@
- Xu et al. (2024) Yao Xu, Shizhu He, Jiabei Chen, Zihao Wang, Yangqiu Song, Hanghang Tong, Guang Liu, Kang Liu, and Jun Zhao. 2024. Generate-on-graph: Treat llm as both agent and kg in incomplete knowledge graph question answering. arXiv preprint arXiv:2404.14741 (2024).
@@
- Xue et al. (2024) Shangzi Xue, Zhenya Huang, Jiayu Liu, Xin Lin, Yuting Ning, Binbin Jin, Xin Li, and Qi Liu. 2024. Decompose, analyze and rethink: Solving intricate problems with human-like reasoning cycle. Advances in Neural Information Processing Systems 37 (2024), 357–385.
@@
- Yih et al. (2016) Wen-tau Yih, Matthew Richardson, Christopher Meek, Ming-Wei Chang, and Jina Suh. 2016. The value of semantic parse labeling for knowledge base question answering. In Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers). 201–206.
@@
- Zeng et al. (2024) Jingying Zeng, Richard Huang, Waleed Malik, Langxuan Yin, Bojan Babic, Danny Shacham, Xiao Yan, Jaewon Yang, and Qi He. 2024. Large language models for social networks: Applications, challenges, and solutions. arXiv preprint arXiv:2401.02575 (2024).
@@
- Zhang et al. (2025) Zhiqiang Zhang, Liqiang Wen, and Wen Zhao. 2025. Rule-KBQA: rule-guided reasoning for complex knowledge base question answering with large language models. In Proceedings of the 31st International Conference on Computational Linguistics. 8399–8417.
@@
- Zhao et al. (2024) Qian Zhao, Hao Qian, Ziqi Liu, Gong-Duo Zhang, and Lihong Gu. 2024. Breaking the barrier: utilizing large language models for industrial recommendation systems through an inferential knowledge graph. In Proceedings of the 33rd ACM International Conference on Information and Knowledge Management. 5086–5093.
@@ A.2. Baselines
We assess the performance of Graph-RFT, leveraging Qwen2.5-7B as the backbone model. The baselines we compare can be divided into three groups: (1) LLM-only methods, including standard In-Context (IO) prompting, Chain-of-Thought (CoT) prompting, and Self-Consistency (SC) prompting. All comparisons utilize six in-content examples and incorporate ”step-by-step” reasoning chains. (2) Semantic Parsing methods, including KB-BINDER(Li et al., 2023a), ChatKBQA(Luo et al., 2023b). (3) Retrieval Augmented methods, including StructGPT(Jiang et al., 2023), RoG(Luo et al., 2023a), ToG(Sun et al., 2023), PoG(Tan et al., 2025), GoG(Xu et al., 2024).
@@ A.3. Implementation Details
We use Qwen2.5-7B-instruct and Qwen2.5-7B-base as backbones (Team et al., 2024) to obtain Graph-RFT-instruct and Graph-RFT-base, respectively, using PyTorch on 8 NVIDIA A800 (80GB) GPUs. In the SFT stage, we leverage LLaMA-Factory framework for efficient LLM fine-tuning. We employ a batchsize of 256 for 2 epochs with a learning rate of 1.4e-5 and AdamW optimizer with cosine decay. In the RL stage, we adopt the Verl framework with the batch size of 128 while the mini-batch size of 32, a rollout number of 8.

### v3

@@ Acknowledgements.
- Deepseekmath: pushing the limits of mathematical reasoning in open language models. arXiv preprint arXiv:2402.03300. Cited by: §3.3.2, §3.3.2.
@@
- Search and refine during think: autonomous retrieval-augmented reasoning of llms. arXiv preprint arXiv:2505.11277. Cited by: §1, §3.3.2.
@@
- Think-on-graph: deep and responsible reasoning of large language model on knowledge graph. arXiv preprint arXiv:2307.07697. Cited by: §A.2, §1, §2, Table 2, Table 2, Table 2, Table 2, Table 2, §4.
@@
- The web as a knowledge-base for answering complex questions. arXiv preprint arXiv:1803.06643. Cited by: §4.
- Reason-rft: reinforcement fine-tuning for visual reasoning of vision language models. In The Thirty-ninth Annual Conference on Neural Information Processing Systems, Cited by: §3.2.
@@
- Paths-over-graph: knowledge graph empowered large language model reasoning. In Proceedings of the ACM on Web Conference 2025, pp. 3505–3522. Cited by: §A.2, Table 1, §1, §1, §2, Table 2.
- Qwen2 technical report. arXiv preprint arXiv:2407.10671 2, pp. 3. Cited by: §A.3.
- Pre-trained language models in biomedical domain: a systematic survey. ACM Computing Surveys 56 (3), pp. 1–52. Cited by: §1.
@@
- Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems 35, pp. 24824–24837. Cited by: §1, §2.
- Medical foundation large language models for comprehensive text analysis and beyond. npj Digital Medicine 8 (1), pp. 141. Cited by: §1.
- Finben: a holistic financial benchmark for large language models. Advances in Neural Information Processing Systems 37, pp. 95716–95743. Cited by: §1.
- Pixiu: a comprehensive benchmark, instruction dataset and large language model for finance. Advances in Neural Information Processing Systems 36, pp. 33469–33484. Cited by: §1.
@@
- Explicit semantic ranking for academic search via knowledge graph embedding. In Proceedings of the 26th international conference on world wide web, pp. 1271–1279. Cited by: §2.
@@
- Generate-on-graph: treat llm as both agent and kg in incomplete knowledge graph question answering. arXiv preprint arXiv:2404.14741. Cited by: §A.2, §1, §2, Table 2, Table 2, §4.
@@
- Decompose, analyze and rethink: solving intricate problems with human-like reasoning cycle. Advances in Neural Information Processing Systems 37, pp. 357–385. Cited by: §1.
@@
- The value of semantic parse labeling for knowledge base question answering. In Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), pp. 201–206. Cited by: §4.
@@
- Large language models for social networks: applications, challenges, and solutions. arXiv preprint arXiv:2401.02575. Cited by: §1.
@@
- Rule-kbqa: rule-guided reasoning for complex knowledge base question answering with large language models. In Proceedings of the 31st International Conference on Computational Linguistics, pp. 8399–8417. Cited by: Table 1, §2.
@@
- Breaking the barrier: utilizing large language models for industrial recommendation systems through an inferential knowledge graph. In Proceedings of the 33rd ACM International Conference on Information and Knowledge Management, pp. 5086–5093. Cited by: §1.
@@ A.2. Baselines
We assess the performance of Graph-RFT, leveraging Qwen2.5-7B as the backbone model. The baselines we compare can be divided into three groups: (1) LLM-only methods, including standard In-Context (IO) prompting, Chain-of-Thought (CoT) prompting, and Self-Consistency (SC) prompting. All comparisons utilize six in-content examples and incorporate ”step-by-step” reasoning chains. (2) Semantic Parsing methods, including KB-BINDER(Li et al., 2023a), ChatKBQA(Luo et al., 2023a). (3) Retrieval Augmented methods, including StructGPT(Jiang et al., 2023), RoG(Luo et al., 2023b), ToG(Sun et al., 2023), PoG(Tan et al., 2025b), GoG(Xu et al., 2024).
@@ A.3. Implementation Details
We use Qwen2.5-7B-instruct and Qwen2.5-7B-base as backbones (Team and others, 2024) to obtain Graph-RFT-instruct and Graph-RFT-base, respectively, using PyTorch on 8 NVIDIA A800 (80GB) GPUs. In the SFT stage, we leverage LLaMA-Factory framework for efficient LLM fine-tuning. We employ a batchsize of 256 for 2 epochs with a learning rate of 1.4e-5 and AdamW optimizer with cosine decay. In the RL stage, we adopt the Verl framework with the batch size of 128 while the mini-batch size of 32, a rollout number of 8.